# Rectified Diffusion Guidance for Conditional Generation

### v1

@@ 3.2 Misconceptions about CFG on Expectation Shift
We first generalize the counterexample in Bradley & Nakkiran (2024) to confirm the expectation shift phenomenon. For VE-SDE with deterministic sampling recipe, we consider the 1-dimensional distribution with $q_{0}(\mathbf{x}_{0}|c)\sim\mathcal{N}(c,1)$, $q(c)\sim\mathcal{N}(0,1)$, $q_{0}(\mathbf{x}_{0})\sim\mathcal{N}(0,2)$. Then we can formulate the forward process and score functions as below:
$q_{t}(\mathbf{x}_{t}|c)$ $\sim\mathcal{N}(c,1+t),$ (18)
$q_{t}(\mathbf{x}_{t})$ $\sim\mathcal{N}(0,2+t),$ (19)
$\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t}|c)$ $=-\frac{\mathbf{x}_{t}-c}{1+t},$ (20)
$\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t})$ $=-\frac{\mathbf{x}_{t}}{2+t}.$ (21)
@@
Denote by $q_{0,\gamma}^{\mathrm{deter}}(\mathbf{x}_{0}|c)$ the conditional distribution by solving PF-ODE in Eq. 10 with CFG weight $\gamma>1$. Then $q_{0,\gamma}^{\mathrm{deter}}(\mathbf{x}_{0}|c)$ follows the closed-form expression as below.
$q_{0,\gamma}^{\mathrm{deter}}(\mathbf{x}_{0}|c)\sim\mathcal{N}\left(c\phi(\gamma,T),2^{1-\gamma}\frac{T+1}{(T+1)^{\gamma}(T+2)^{1-\gamma}}\right),$ (22)
@@
$\phi(\gamma,T)=2^{\frac{1-\gamma}{2}}\left(\frac{1}{(T+1)^{\frac{\gamma}{2}}(T+2)^{\frac{1-\gamma}{2}}}+\frac{\gamma}{2}\int_{0}^{T}(s+1)^{-\frac{\gamma+2}{2}}(s+2)^{-\frac{1-\gamma}{2}}\mathrm{d}s\right).$ (23)
@@
Furthermore, we have closed-form expression for $\phi(\gamma)$ when $\gamma\in\mathbb{N}$ and $\gamma>1$, i.e.,
$\phi(\gamma)=\begin{cases}{2^{-n}\left(\sum_{k=0}^{n}C_{n}^{k}\frac{2n+1}{2n-2k+1}\right)},&\gamma=2n+1,\\[20.0pt] {2^{-\frac{1}{2}}\left(\sqrt{2}-\frac{1}{2}\log\frac{\sqrt{2}-1}{\sqrt{2}+1}\right)},&\gamma=2,\\[20.0pt] {\frac{(2n-1)!!\sqrt{2}n}{(2n)!!2^{n}}\left(\left(\sum_{k=2}^{n}\frac{1}{k}\frac{(2k)!!2^{k-\frac{1}{2}}}{(2k-1)!!}\right)+2\sqrt{2}-\log\frac{\sqrt{2}-1}{\sqrt{2}+1}\right)},&\gamma=2n\geqslant 4.\end{cases}$ (26)
@@ 3.3 Rectified Classifier-Free Guidance
$\nabla_{\mathbf{x}_{t}}\log q_{t,\gamma_{1},\gamma_{0}}(\mathbf{x}_{t}|c)=\gamma_{1}\otimes\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t}|c)+\gamma_{0}\otimes\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t}),$ (27)
in which $\gamma_{1}$, $\gamma_{0}\in\mathbb{R}^{D}$ are functions with respect to condition $c$ and timestep $t$, and $\otimes$ indicates element-wise product. Denote by $q_{0,\gamma_{1},\gamma_{0}}(\mathbf{x}_{0}|c)$ the attached conditional distribution following PF-ODE in Eq. 10 with $\nabla_{\mathbf{x}_{t}}\log q_{t,\gamma_{1},\gamma_{0}}(\mathbf{x}_{t}|c)$.
@@
  • Each component $\gamma_{1,i}>1$ for strengthened conditional fidelity,
  • Denoising with PF-ODE and Eq. 27 is theoretically the reciprocal of forward process, thus $q_{0,\gamma_{1},\gamma_{0}}(\mathbf{x}_{0}|c)$ enjoys the same expectation as ground-truth $q_{0}(\mathbf{x}_{0}|c)$,
  • $q_{0,\gamma_{1},\gamma_{0}}(\mathbf{x}_{0}|c)$ enjoys smaller or the same variance as ground-truth $q_{0}(\mathbf{x}_{0}|c)$ for sharper distribution and better exemplars.
@@
$\mathbb{E}_{q_{t}(\mathbf{x}_{t}|c)}[\nabla_{\mathbf{x}_{t}}\log q_{t,\gamma_{1},\gamma_{0}}(\mathbf{x}_{t}|c)]=-\frac{1}{\sigma_{t}}\mathbb{E}_{q_{t}(\mathbf{x}_{t}|c)}[\gamma_{1}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},c,t)+\gamma_{0}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)]=0.$ (28)
To confirm the feasibility and precisely describe the expectation of $q_{0,\gamma_{1},\gamma_{0}}(\mathbf{x}_{0}|c)$, resembling Eqs. 14 and 15 we can write:
$\mathbf{f}_{\theta,\gamma_{1},\gamma_{0}}^{t}(\mathbf{x}_{t},c)$ $=\frac{1}{\alpha_{t}}(\mathbf{x}_{t}-\sigma_{t}(\gamma_{1}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},c,t)+\gamma_{0}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t))),$ (29)
$p_{\theta,\gamma_{1},\gamma_{0}}(\mathbf{x}_{t-1}|\mathbf{x}_{t},c)$ =$=\begin{cases}{q_{\delta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{f}_{\theta,\gamma_{1},\gamma_{0}}^{t}(\mathbf{x}_{t},c),c)},&t>1,\\[5.0pt] {\mathcal{N}(\mathbf{f}_{\theta,\gamma_{1},\gamma_{0}}^{t}(\mathbf{x}_{1},c),\sigma_{1}^{2}\mathbf{I})},&t=1.\end{cases}$ (30)
And we have the following theorem, in which proof is addressed in Section A.3.
@@
Let $\mathbf{x}_{t}\sim q_{t}(\mathbf{x}_{t}|c)$, $\tilde{\mathbf{x}}_{t}\sim p_{\theta,\gamma_{1},\gamma_{0}}(\tilde{\mathbf{x}}_{t}|c)$ induced from DDIM sampler in Eq. 30. Assume that all $\delta_{t}=0$, denote by $\Delta_{t}$ the difference between expectation of $\mathbf{x}_{t}$ and $\tilde{\mathbf{x}}_{t}$, i.e.,
$\Delta_{t}=\mathbb{E}_{q_{t}(\mathbf{x}_{t}|c)}[\mathbf{x}_{t}]-\mathbb{E}_{p_{\theta,\gamma_{1},\gamma_{0}}(\tilde{\mathbf{x}}_{t}|c)}[\tilde{\mathbf{x}}_{t}].$ (31)
@@
$\Delta_{t-1}=\frac{\sigma_{t-1}}{\sigma_{t}}\Delta_{t}-(\sigma_{t-1}-\frac{\alpha_{t-1}}{\alpha_{t}}\sigma_{t})\mathbb{E}_{\tilde{\mathbf{x}}_{t}}[(\gamma_{1}-1)\bm{\epsilon}_{\theta}(\tilde{\mathbf{x}}_{t},c,t)+\gamma_{0}\bm{\epsilon}_{\theta}(\tilde{\mathbf{x}}_{t},t)]).$ (32)
@@
$\Delta_{t-1}$ $=-(\sigma_{t-1}-\frac{\alpha_{t-1}}{\alpha_{t}}\sigma_{t})\mathbb{E}_{\mathbf{x}_{t}}[(\gamma_{1}-1)\bm{\epsilon}_{\theta}(\mathbf{x}_{t},c,t)+\gamma_{0}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)]).$ (33)
Theorem 3 studies the difference between expectation of denoising with Eq. 27 and the ground-truth. Note that $\mathbb{E}_{\mathbf{x}_{t}}[\bm{\epsilon}_{\theta}(\mathbf{x}_{t},c,t)]=\mathbb{E}_{\mathbf{x}_{t}}[\mathbb{E}_{q_{t}(\bm{\epsilon}|\mathbf{x}_{t})}[\bm{\epsilon}|\mathbf{x}_{t}]]=\mathbb{E}_{\mathbf{x}_{t}}[\bm{\epsilon}]=0$, we have
$\mathbb{E}_{\mathbf{x}_{t}}[(\gamma_{1}-1)\bm{\epsilon}_{\theta}(\mathbf{x}_{t},c,t)+\gamma_{0}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)]=\mathbb{E}_{\mathbf{x}_{t}}[\gamma_{1}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},c,t)+\gamma_{0}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)],$ (34)
which coincides with Eq. 28, indicating the feasibility and a closed-form solution given $c$ and $t$.
As for variance, however, normally we cannot calculate the variance of $p_{\theta,\gamma_{1},\gamma_{0}}(\mathbf{x}_{t}|c)$. Instead, we study the variance in Sec. 3.2 as an empirical evidence, where proof is in Section A.4.
@@
Under settings in Theorem 2, denote by $q_{0,\gamma_{1},\gamma_{0}}^{\mathrm{deter}}(\mathbf{x}_{0}|c)$ the conditional distribution by PF-ODE and deterministic sampling with $\gamma_{1}$ and $\gamma_{0}$ as in Eq. 27. Then we have
@@
According to Theorem 4, variance of $q_{0,\gamma_{1},\gamma_{0}}^{\mathrm{deter}}(\mathbf{x}_{0}|c)$ is guaranteed to be smaller than the ground-truth $\mathrm{var}_{q_{0}(\mathbf{x}_{0}|c)}[\mathbf{x}_{0}]=1$ when each $\gamma_{0,i}\leqslant 0$ and $\gamma_{1,i}+\gamma_{0,i}\geqslant 1$, especially when $T\rightarrow+\infty$.
Now we formally propose the constraints. First, we need $\gamma_{1,i}>1$ for strengthened conditional fidelity. Then for expectation, it is noteworthy that $\Delta_{T}=0$ satisfies the assumption in Theorem 3. Therefore by induction, it is feasible to annihilate $\Delta_{0}$ by annihilation of Eq. 28. Finally as for variance, we empirically set $\gamma_{0,i}\leqslant 0$ and $\gamma_{1,i}+\gamma_{0,i}\geqslant 0$.
Practically, we can determine $\gamma_{0}$ according to the guidance strength $\gamma_{1}$, condition $c$, and timestep $t$, according to the closed-form solution of Eq. 33. Concretely, given condition $c$, it is feasible to pre-compute a collection of $\{(\epsilon_{\theta}(\mathbf{x}_{t},c,t),\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t))\}_{t}$ by traversing $q_{0}(\mathbf{x}_{0}|c)$, and maintain a lookup table consisting of $\mathbb{E}_{\mathbf{x}_{t}}[\bm{\epsilon}_{\theta}(\mathbf{x}_{t},c,t)]/\mathbb{E}_{\mathbf{x}_{t}}[\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)]$. Then given any $\gamma_{1}$, we can directly achieve $\gamma_{0}$ by multiplying $-(\gamma_{1}-1)$ with the expectation ratio. Pseudo-code is addressed in Appendix B.
@@
$\|\Delta_{t-1}\|_{2}^{2}$ $\leqslant(\sigma_{t-1}-\frac{\alpha_{t-1}}{\alpha_{t}}\sigma_{t})^{2}\mathbb{E}_{\mathbf{x}_{t}}[\|(\gamma_{1}-1)\bm{\epsilon}_{\theta}(\mathbf{x}_{t},c,t)+\gamma_{0}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)\|_{2}^{2}].$ (36)
Then we can define the objective resembling DPMs as below, optimizing reversely from $t=T$ to 00.
$\mathcal{L}_{\gamma_{1},\gamma_{0}}=\mathbb{E}_{\mathbf{x}_{t},t}[\|(\gamma_{1}-1)\bm{\epsilon}_{\theta}(\mathbf{x}_{t},c,t)+\gamma_{0}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)\|_{2}^{2}].$ (37)
@@
For any $\{\delta_{t}\}_{t}$, ReCFG with $\mathcal{L}_{\gamma_{1},\gamma_{0}}$ is compatible with native DPM up to a constant.

### v2

@@ 3.2 Misconceptions on Expectation Shift
We first generalize the counterexample in [1] to confirm the expectation shift phenomenon. For VE-SDE with deterministic sampling recipe, we consider the 1-dimensional distribution with $q_{0}(\mathbf{x}_{0}|c)\sim\mathcal{N}(c,1)$, $q(c)\sim\mathcal{N}(0,1)$, $q_{0}(\mathbf{x}_{0})\sim\mathcal{N}(0,2)$. Then we can formulate the forward process and score functions as below:
$\hskip-11.38109ptq_{t}(\mathbf{x}_{t}|c)\sim\mathcal{N}(c,1+t),$ $\;\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t}|c)=-\frac{\mathbf{x}_{t}-c}{1+t},$ (20)
$\hskip-11.38109ptq_{t}(\mathbf{x}_{t})\sim\mathcal{N}(0,2+t),$ $\;\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t})=-\frac{\mathbf{x}_{t}}{2+t}.$ (21)
@@
Denote by $q_{0,\gamma}^{\mathrm{deter}}(\mathbf{x}_{0}|c)$ the conditional distribution by solving PF-ODE in Eq. 10 with $\gamma>1$. Then $q_{0,\gamma}^{\mathrm{deter}}(\mathbf{x}_{0}|c)$ follows the closed-form expression as below.
$q_{0,\gamma}^{\mathrm{deter}}(\mathbf{x}_{0}|c)\sim\mathcal{N}\left(c\phi(\gamma,T),2^{1-\gamma}\psi(\gamma,T)\right),$ (22)
@@
$\phi(\gamma,T)$ $=\frac{2^{\frac{1-\gamma}{2}}}{(T+1)^{\frac{\gamma}{2}}(T+2)^{\frac{1-\gamma}{2}}}$ (23)
$\qquad+\frac{\gamma}{2^{\frac{\gamma+1}{2}}}\int_{0}^{T}\frac{(s+1)^{-\frac{\gamma+2}{2}}}{(s+2)^{\frac{1-\gamma}{2}}}\mathrm{d}s,$ (24)
$\psi(\gamma,T)$ $=\frac{T+1}{(T+1)^{\gamma}(T+2)^{1-\gamma}}.$ (25)
Figure 2: Visualization of the lookup table on LDM [25], EDM2 [16], and SD3 [9], each of which consists of the expectation ratio $\mathbb{E}_{\mathbf{x}_{t}}[\bm{\epsilon}_{\theta}(\mathbf{x}_{t},c,t)]/\mathbb{E}_{\mathbf{x}_{t}}[\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)]$. Each pixel represents the scale of the pixel-wise ratio, i.e., color red implies that ratio is greater than one, while color blue stands for ratio smaller than one. The darker the color is, the farther the ratio appears away from one. We report in each row the expectation ratios on five timesteps uniformly sampled from the whole trajectory, under different DPMs and NFEs. It is noteworthy that expectation ratios at the same timestep vary largely by different pixels, and there is no general pattern along with timesteps or pixels.
@@ 3.3 Rectified Classifier-Free Guidance
$s_{t,\gamma_{1},\gamma_{0}}(\mathbf{x},c)$ $=\gamma_{1}\otimes\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x}|c)$ (28)
$\qquad+\gamma_{0}\otimes\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x}),$ (29)
in which $\gamma_{1}$, $\gamma_{0}\in\mathbb{R}^{D}$ are functions with respect to condition $c$ and timestep $t$, and $\otimes$ indicates element-wise product. Denote by $q_{0,\gamma_{1},\gamma_{0}}^{\mathrm{deter}}(\mathbf{x}_{0}|c)$ the attached conditional distribution following PF-ODE in Eq. 10 with $s_{t,\gamma_{1},\gamma_{0}}(\mathbf{x},c)$.
@@
  • Each component of $\gamma$ is larger than one for strengthened conditional fidelity, i.e., $\gamma_{1,i}>1$,
  • Denoising with PF-ODE and Eq. 28 is theoretically the reciprocal of forward process, thus $q_{0,\gamma_{1},\gamma_{0}}^{\mathrm{deter}}(\mathbf{x}_{0}|c)$ enjoys the same expectation as the ground-truth $q_{0}(\mathbf{x}_{0}|c)$,
  • $q_{0,\gamma_{1},\gamma_{0}}^{\mathrm{deter}}(\mathbf{x}_{0}|c)$ enjoys smaller or the same variance as the ground-truth $q_{0}(\mathbf{x}_{0}|c)$ for sharper distribution and thus concentrated better exemplars.
@@
$\mathbb{E}_{q_{t}(\mathbf{x}_{t}|c)}[s_{t,\gamma_{1},\gamma_{0}}(\mathbf{x},c)]=\mathbf{0}.$ (30)
To confirm the feasibility and precisely describe the expectation of $q_{0,\gamma_{1},\gamma_{0}}^{\mathrm{deter}}(\mathbf{x}_{0}|c)$, resembling Eqs. 16 and 17 we write denoised observation and denoising process as below:
$\hat{\bm{\epsilon}}_{\theta}(\mathbf{x}_{t},c,t)$ $=\gamma_{1}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},c,t)+\gamma_{0}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t),$ (31)
$\hat{\mathbf{f}}_{\theta}^{t}(\mathbf{x}_{t},c)$ $=\frac{1}{\alpha_{t}}(\mathbf{x}_{t}-\sigma_{t}\hat{\bm{\epsilon}}_{\theta}(\mathbf{x}_{t},c,t)),$ (32)
$\hskip-5.69054pt\hat{p}_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},c)$ $=\begin{cases}{q_{\delta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\hat{\mathbf{f}}_{\theta}^{t}(\mathbf{x}_{t},c),c)},&t>1,\\[2.0pt] {\mathcal{N}(\hat{\mathbf{f}}_{\theta}^{t}(\mathbf{x}_{1},c),\sigma_{1}^{2}\mathbf{I})},&t=1.\end{cases}$ (33)
We have the theorem below, proof is in Section A.3.
@@
Let $\mathbf{x}_{t}\sim q_{t}(\mathbf{x}_{t}|c)$, $\tilde{\mathbf{x}}_{t}\sim\hat{p}_{\theta}(\tilde{\mathbf{x}}_{t}|c)$ induced from DDIM sampler in Eq. 33. Assume that all $\delta_{t}=0$, denote by $\Delta_{t}$ the difference between expectation of $\mathbf{x}_{t}$ and $\tilde{\mathbf{x}}_{t}$, by $\bm{\epsilon}_{\gamma_{1},\gamma_{0}}^{c,t}$ the interpolation between score functions, i.e.,
$\Delta_{t}$ $=\mathbb{E}_{q_{t}(\mathbf{x}_{t}|c)}[\mathbf{x}_{t}]-\mathbb{E}_{\hat{p}_{\theta}(\tilde{\mathbf{x}}_{t}|c)}[\tilde{\mathbf{x}}_{t}],$ (34)
$\bm{\epsilon}_{\gamma_{1},\gamma_{0}}^{c,t}(\mathbf{x})$ $=(\gamma_{1}-1)\bm{\epsilon}_{\theta}(\mathbf{x},c,t)+\gamma_{0}\bm{\epsilon}_{\theta}(\mathbf{x},t).$ (35)
@@
$\hskip-8.53581pt\Delta_{t-1}=\frac{\sigma_{t-1}}{\sigma_{t}}\Delta_{t}-(\sigma_{t-1}-\frac{\alpha_{t-1}}{\alpha_{t}}\sigma_{t})\mathbb{E}_{\tilde{\mathbf{x}}_{t}}[\bm{\epsilon}_{\gamma_{1},\gamma_{0}}^{c,t}(\tilde{\mathbf{x}}_{t})].$ (36)
@@
$\Delta_{t-1}$ $=-(\sigma_{t-1}-\frac{\alpha_{t-1}}{\alpha_{t}}\sigma_{t})\mathbb{E}_{\mathbf{x}_{t}}[\bm{\epsilon}_{\gamma_{1},\gamma_{0}}^{c,t}(\mathbf{x}_{t})].$ (37)
Theorem 3 studies the difference between expectation of denoising with Eq. 28 and the ground-truth. Note that
$\mathbb{E}_{\mathbf{x}_{t}}[\bm{\epsilon}_{\theta}(\mathbf{x}_{t},c,t)]=\mathbb{E}_{\mathbf{x}_{t}}[\mathbb{E}[\bm{\epsilon}|\mathbf{x}_{t}]]=\mathbb{E}_{\mathbf{x}_{t}}[\bm{\epsilon}]=\mathbf{0},$ (38)
therefore we have
$\;\mathbb{E}_{\mathbf{x}_{t}}[\bm{\epsilon}_{\gamma_{1},\gamma_{0}}^{c,t}(\mathbf{x}_{t})]$ (39)
$=$ $\;\mathbb{E}_{\mathbf{x}_{t}}[(\gamma_{1}-1)\bm{\epsilon}_{\theta}(\mathbf{x}_{t},c,t)+\gamma_{0}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)]$ (40)
$=$ $\;\mathbb{E}_{\mathbf{x}_{t}}[\gamma_{1}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},c,t)+\gamma_{0}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)],$ (41)
which coincides with Eq. 30, indicating the feasibility and a closed-form solution given $c$ and $t$ as below:
$\gamma_{0}=(1-\gamma_{1})\mathbb{E}_{\mathbf{x}_{t}}[\bm{\epsilon}_{\theta}(\mathbf{x}_{t},c,t)]/\mathbb{E}_{\mathbf{x}_{t}}[\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)].$ (42)
As for variance, however, normally we cannot analytically calculate the variance of $\hat{p}_{\theta}(\mathbf{x}_{t}|c)$. Instead, we study the variance of toy data in Sec. 3.2 as an empirical evidence in the following theorem, where proof is in Section A.4.
@@
Under settings in Theorem 2, denote by $q_{0,\gamma_{1},\gamma_{0}}^{\mathrm{deter}}(\mathbf{x}_{0}|c)$ the conditional distribution by PF-ODE with $\gamma_{1}$ and $\gamma_{0}$ as in Eq. 28. Then we have
@@
Figure 3: Qualitative comparison on EDM2 and SD3. Left and right in each cell suggest samples via CFG and ReCFG, respectively.
According to Theorem 4, it is noteworthy that variance of $q_{0,\gamma_{1},\gamma_{0}}^{\mathrm{deter}}(\mathbf{x}_{0}|c)$ under toy setting is guaranteed to be smaller than the ground-truth $\mathrm{var}_{q_{0}(\mathbf{x}_{0}|c)}[\mathbf{x}_{0}]=1$ when each component satisfies that $\gamma_{0,i}\leqslant 0$ and $\gamma_{1,i}+\gamma_{0,i}\geqslant 1$, especially when $T\rightarrow+\infty$.
Now we formally propose the constraints. First, we need each component $\gamma_{1,i}>1$ for strengthened conditional fidelity. Then for expectation, it is noteworthy that $\Delta_{T}=0$ satisfies the assumption in Theorem 3. Therefore by induction, it is feasible to annihilate $\Delta_{0}$ by annihilation of Eq. 30 at all intermediate timesteps $t$. Finally as for variance, we empirically set $\gamma_{0,i}\leqslant 0$ and $\gamma_{1,i}+\gamma_{0,i}\geqslant 0$.
Practically, we can determine $\gamma_{0}$ according to the guidance strength $\gamma_{1}$, condition $c$, and timestep $t$, according to the closed-form solution in Eq. 42. Concretely, given condition $c$, it is feasible to pre-compute a collection of $\{(\epsilon_{\theta}(\mathbf{x}_{t},c,t),\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t))\}_{t}$ by traversing $q_{0}(\mathbf{x}_{0}|c)$, and maintain a lookup table consisting of $\mathbb{E}_{\mathbf{x}_{t}}[\bm{\epsilon}_{\theta}(\mathbf{x}_{t},c,t)]/\mathbb{E}_{\mathbf{x}_{t}}[\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)]$. Then given any $\gamma_{1}$, we can directly achieve $\gamma_{0}$ by multiplying $(\gamma_{1}-1)$ with the expectation ratio. Pseudo-code is addressed in Appendix B.
@@
$\|\Delta_{t-1}\|_{2}^{2}$ $\leqslant(\sigma_{t-1}-\frac{\alpha_{t-1}}{\alpha_{t}}\sigma_{t})^{2}\mathbb{E}_{\mathbf{x}_{t}}[\|\bm{\epsilon}_{\gamma_{1},\gamma_{0}}^{c,t}(\mathbf{x}_{t})\|_{2}^{2}].$ (44)
Then we can define the objective resembling DPMs as below, optimizing reversely from $t=T$ to 0.
$\mathcal{L}=\mathbb{E}_{\mathbf{x}_{t},t}[\|(\gamma_{1}-1)\bm{\epsilon}_{\theta}(\mathbf{x}_{t},c,t)+\gamma_{0}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)\|_{2}^{2}].$ (45)
@@
For any $\{\delta_{t}\}_{t}$, ReCFG with $\mathcal{L}$ is compatible with native DPM up to a constant.